boredfr:

The length of the side of a square is represented by the expression 2𝑎^2b. Which term represents the area of this square? A. 8a^3b B. 4a^3b C. 4a^4b^2 D. 2a^6b^2 @angel

boredfr:

@Angle *****

Angle:

What is general formula for the area of a square?

boredfr:

Side times side

Angle:

ok, so if a side in this problem is ` 2 a^2 b ` Then the area = \((2a^2b)(2a^2b)\)

Angle:

we want to simplify that

Angle:

do you have any ideas on how to simplify expressions like that?

boredfr:

Yeah that would be 2*2=4 a^2 * a^2 = a^4 and b*b=b^2 4a^4b^2 ?

Angle:

perfect!! :)

boredfr:

Thank you ;p